# Université de Genève, CH-1211 Genève, Switzerland

### v1

@@ 3.3 The wave function in the rational case, I
$\widetilde{\varphi}^{(+)}_{k}(X_{D})$ $=\sum_{s\geq 0}{\hat{a}_{sP+k}(q_{D})\over(sP+k)q_{D}^{\sigma_{sP+k}}}X_{D}^{sP+k}\,,\quad k=1,\dots,P-1\,,$ (3.78c)
$\widetilde{\varphi}^{(-)}_{k}(X_{D})$ $=\sum_{s\geq 0}{\hat{a}_{-(sP+k)}(q_{D})\over-(sP+k)q_{D}^{sP+k+\sigma_{-(sP+k)}}}X_{D}^{-(sP+k)}\,,\quad k=1,\dots,P-1\,,$ (3.78d)
@@
$\widetilde{\varphi}_{k}(X_{D})$ $=\widetilde{\varphi}^{(+)}_{k}(X_{D})-q_{D}^{k}\widetilde{\varphi}^{(-)}_{k}(X_{D})\,,\quad k=1,\dots,Q-1\,.$ (3.79b)
Putting Eqs. (3.76), (3.77) together and suitably identifying the functions just introduced in Eq. (3.3), we can now write the antiholomorphic block as
@@
$\widetilde{\varphi}^{(+)}_{P}(X_{D})={Q\over P}\varphi^{(+)}_{Q}(X)\,,\quad\widetilde{\varphi}^{(-)}_{P}(X_{D})={Q\over P}q^{sQ\left(1-\frac{Q}{P}\right)}\varphi^{(-)}_{Q}(X)\,,$ (3.81)
as a direct consequence of the cancellation symmetry in Eq. (3.61). Therefore, after summing Eqs. (3.73), (3.80) and using the relations in Eq. (3.81), we obtain the expression
$\phi^{\rm WKB}(X;q)+\phi^{\rm WKB}(X_{D};q_{D})=$ $\left(-{1\over 2}-{Q\over 2P}-{{\rm i}Q\over 2\pi P}\log X\right)\varphi_{Q}(X)$ (3.82)
$+\left({{\rm i}Q\over 2\pi P}+2q{\partial\over\partial q}\right)\int_{0}^{X}\varphi_{Q}(X^{\prime}){{\rm d}X^{\prime}\over X^{\prime}}$
$+sQ\left(1-\frac{Q}{P}\right)\int_{0}^{X}\varphi^{(-)}_{Q}(X^{\prime}){{\rm d}X^{\prime}\over X^{\prime}}$
$+\sum_{k=1}^{Q-1}{\varphi_{k}(X)\over q^{k}-1}+\sum_{k=1}^{P-1}{\widetilde{\varphi}_{k}(X^{Q\over P})\over q^{k{Q^{2}\over P^{2}}}-1}\,.$
Moreover, the power series defined in Eqs. (3.3), (3.3), which dictate the formula for the exact wave function at rational points following Eq. (3.82), are determined by the function
@@
$\varphi^{(\pm)}_{k}(X)$ $={1\over Q}\sum_{m=0}^{Q-1}q^{\mp mk}\,\Upsilon^{(\pm)}(q^{m}X,q)\,,\quad k=1,\dots,Q-1\,,$ (3.84b)
@@
$\widetilde{\varphi}^{(\pm)}_{k}(X_{D})$ $={1\over P}\sum_{m=0}^{P-1}q_{D}^{\mp mk}\,\Upsilon^{(\pm)}(q_{D}^{m}X_{D})\,,\quad k=1,\dots,P-1\,.$ (3.85b)
Let us stress that the derivation we have presented here implies a fixed choice of the classical branch. Indeed, the functions $\Upsilon(X),\varphi_{Q}(X),\varphi_{k}(X),\widetilde{\varphi}_{P}(X_{D}),\widetilde{\varphi}_{k}(X_{D})$ are implicitly dependent on the label $\alpha$. We note that, applying Eqs. (3.65), (3.66), we can simply read off the integer coefficients $c_{\pm s}$, $s\in{\mathbb{Z}}_{>0}$, from the function $\varphi_{Q}(X)$. Namely, we have that
@@
after substituting $q^{sQ}=1$. In Appendix B, we describe an alternative way of computing the integer sequences $\{c_{\pm s}\}$, $s\in{\mathbb{Z}}_{>0}$ using Eq. (3.82). We will show in Section 4 that, starting from the known closed expression for the holomorphic and antiholomorphic blocks, we can successfully compute the integer coefficients for both series in $X,X^{-1}$ and all choices of the saddle connection $\alpha$ in the examples of the figure-eight and the three-twist knots.
@@ 3.4 The wave function in the rational case, II
In Section 3.3, we showed that the exact wave function $\chi(u;\hbar)$ in Eq. (3.51) at rational values of $\hbar$ is determined by the functions $\Upsilon^{(\alpha)}(X)$, defined in Eq. (3.83), and their derivatives with respect to $q$, up to the constants $\mathsf{C}_{\alpha}$ and the exponential prefactors containing $s^{(\alpha)}_{0,1}(x)$. We will now prove that the functions $\Upsilon^{(\alpha)}(X)$ can be obtained by solving directly the quantum $A$-polynomial in closed form. We will focus on $q$-difference equations of order two and three. These are the relevant cases for the examples studied in Section 4. The matrix formalism presented below was used in szabolcs for a second-order $q$-difference equation in the context of quantum mirror curves.
3.4.1 Solving a second-order $q$-difference equation
Recall that, for each choice of $\alpha$, the resummed perturbative wave function $\chi^{(\alpha),\rm{WKB}}(u;\hbar)$ is given in Eq. (3.39). Equivalently, we write
$\chi^{(\alpha),\rm{WKB}}(u;\hbar)={\rm e}^{\frac{{\rm i}}{\hbar}s^{(\alpha)}_{0}(x)+s^{(\alpha)}_{1}(x)}\Psi^{(\alpha),\rm{WKB}}(X;q)\,,$ (3.87)
@@ 3.4.1 Solving a second-order $q$-difference equation
$\Psi^{(\alpha),\rm{WKB}}(X;q)=\exp\left(\phi^{(\alpha),\rm{WKB}}(X;q)\right)\,,$ (3.88)
@@
$\hat{A}(\hat{P},\hat{X},q)\chi^{(\alpha),\rm{WKB}}(u;\hbar)=0\,,$ (3.89)
is of second order. After appropriately accounting for the correction due to the exponential prefactor containing the functions $s^{(\alpha)}_{0,1}(x)$ in Eq. (3.87), the $q$-difference equation above becomes
$\Psi^{(\alpha),\rm{WKB}}(qX;q)+A(X)\Psi^{(\alpha),\rm{WKB}}(X;q)+B(X)\Psi^{(\alpha),\rm{WKB}}(q^{-1}X;q)=0\,,$ (3.90)
where $A(X),B(X)$ are functions of $X$ and $q$. Let us observe that, due to the conjectural structure in Eqs. (3.40), (3.41), we have
$\upsilon^{(\alpha)}(X)={\Psi^{(\alpha),\rm{WKB}}(qX;q)\over\Psi^{(\alpha),\rm{WKB}}(X;q)}=\exp\left(\Upsilon^{(\alpha)}(X)\right)\,,$ (3.91)
and we can now write Eq. (3.90) in terms of the above function $\upsilon^{(\alpha)}(X)$ as
$\upsilon(X)\upsilon(q^{-1}X)+A(X)\upsilon(q^{-1}X)+B(X)=0\,,$ (3.92)
where we have dropped the label $\alpha$ from the notation for simplicity. Let us further assume that $\hbar$ is fixed as in Eq. (3.53), that is,
$q={\rm e}^{{\rm i}\hbar}={\rm e}^{2\pi{\rm i}{P\over Q}}\,,\quad P,Q\in{\mathbb{Z}}_{>0}\;\;\text{coprime}\,,$ (3.93)
@@
$\upsilon_{k}=\upsilon(q^{-k}X)\,.$ (3.94)
@@
$\upsilon_{k}\upsilon_{k+1}+A(q^{-k}X)\upsilon_{k+1}+B(q^{-k}X)=0\,,\quad 0\leq k\leq Q-1\,,$ (3.95)
as a consequence of Eq. (3.92). We solve this system recursively, as follows. We define the sequences $\{a^{(k)}(X)\}$ and $\{b^{(k)}(X)\}$, $k\in{\mathbb{Z}}_{\geq 0}$, via the relations
@@
$\upsilon_{0}\upsilon_{1}\cdots\upsilon_{k}+a^{(k)}(X)\upsilon_{k}+b^{(k)}(X)=0\,,\quad k>1\,.$ (3.96c)
If we multiply Eq. (3.96c) by $\upsilon_{k+1}$ and apply Eq. (3.95), we obtain a recursion relation which can be written in matrix form as
$\begin{pmatrix}a^{(k+1)}(X)\\ b^{(k+1)}(X)\end{pmatrix}=\begin{pmatrix}-A(q^{-k}X)&1\\ -B(q^{-k}X)&0\end{pmatrix}\begin{pmatrix}a^{(k)}(X)\\ b^{(k)}(X)\end{pmatrix}\,,$ (3.97)
@@
${\cal M}_{Q}(X)=\prod_{k=1}^{Q}\begin{pmatrix}-A(q^{k}X)&1\\ -B(q^{k}X)&0\end{pmatrix}=\begin{pmatrix}{\cal M}_{11}(X)&{\cal M}_{12}(X)\\ {\cal M}_{21}(X)&{\cal M}_{22}(X)\end{pmatrix}\,,$ (3.99)
@@
${\mathcal{M}}_{Q}(q^{-1}X)$ $=\begin{pmatrix}-A(X)&1\\ -B(X)&0\end{pmatrix}{\mathcal{M}}_{Q}(X)\begin{pmatrix}-A(X)&1\\ -B(X)&0\end{pmatrix}^{-1}$ (3.100)
$=\begin{pmatrix}{\cal M}_{22}-A{\cal M}_{12}&\frac{A}{B}\left({\cal M}_{11}+A{\cal M}_{12}\right)-\frac{1}{B}\left({\cal M}_{21}+A{\cal M}_{22}\right)\\ -B{\cal M}_{12}&{\cal M}_{11}+A{\cal M}_{12}\end{pmatrix}\,,$
where we have removed the explicit dependence on $X$ for compactness. Following Eqs. (3.98), (3.100), we find that [^9] There is a sign misprint in the corresponding equation in szabolcs .
$\begin{pmatrix}a^{(Q)}(q^{-1}X)\\ b^{(Q)}(q^{-1}X)\end{pmatrix}=\begin{pmatrix}{-\mathcal{M}}_{22}(X)+A(X){\mathcal{M}}_{12}(X)\\ B(X){\mathcal{M}}_{12}(X)\end{pmatrix}\,.$ (3.101)
@@
which is invariant under $q$-shifts. We consider Eq. (3.96c) for $k=Q$ and, shifting $X$ into $q^{-1}X$, we produce the system
$\begin{cases}\upsilon_{0}(\Pi\upsilon+a^{(Q)}(X))+b^{(Q)}(X)=0\,,\\ \upsilon_{1}(\Pi\upsilon+a^{(Q)}(q^{-1}X))+b^{(Q)}(q^{-1}X)=0\,.\end{cases}$ (3.103)
By substituting the expressions for $a^{(Q)}(X),a^{(Q)}(q^{-1}X)$ and $b^{(Q)}(X),b^{(Q)}(q^{-1}X)$ in terms of the entries of the original matrix $\mathcal{M}_{Q}(X)$ in Eqs. (3.98), (3.101) and then using Eq. (3.95) with $k=0$, we can rewrite the system above as
$\begin{cases}\upsilon_{0}(\Pi\upsilon-{\mathcal{M}}_{11}(X))-{\mathcal{M}}_{21}(X)=0\,,\\ \Pi\upsilon-{\mathcal{M}}_{22}(X)-\upsilon_{0}{\mathcal{M}}_{12}(X)=0\,,\end{cases}$ (3.104)
@@
$\Pi\upsilon=\frac{{\rm tr}{\mathcal{M}_{Q}}(X)\pm\sqrt{\Delta_{Q}(X)}}{2}\,,$ (3.106)
@@
$\Delta_{Q}(X)=({\rm tr}{\mathcal{M}_{Q}}(X))^{2}-4\,\,{\rm det}{\mathcal{M}_{Q}}(X)\,,$ (3.107)
while ${\rm det}{\mathcal{M}_{Q}}(X)$ and ${\rm tr}{\mathcal{M}_{Q}}(X)$ denote the trace and determinant of the matrix, respectively. Note that, following Eq. (3.99), both ${\rm det}{\mathcal{M}_{Q}}(X)$ and ${\rm tr}{\mathcal{M}_{Q}}(X)$ are invariant under $q$-shifts and so depend on $X$ through $X^{Q}$. The function $\upsilon_{0}$ can be found, for example, using the second line of Eq. (3.104), yielding
$\upsilon_{0}=\frac{\Pi v-{\mathcal{M}}_{22}(X)}{{\mathcal{M}}_{12}(X)}=\frac{{\mathcal{M}}_{11}(X)-{\mathcal{M}}_{22}(X)\pm\sqrt{\Delta_{Q}(X)}}{2{\mathcal{M}}_{12}(X)}\,,$ (3.108)
while the functions $\upsilon_{k}$, $k\in{\mathbb{Z}}_{>0}$, are obtained by successively $q$-shifting $\upsilon_{0}$. This gives the solution to Eq. (3.92) for all rational values of $\hbar$. However, we stress that, for such a solution to be consistent with the original choice of a classical branch labeled by $\alpha$, we must impose that, taking $P=Q=1$, we find
@@
where $P^{(\alpha)}$ is the selected non-abelian solution of the classical $A$-polynomial at fixed $X$ after taking into account the corrections due to the exponential prefactor in Eq. (3.87). In this way, we remove the sign ambiguity in Eq. (3.106) and correspondingly in Eq. (3.108). We conclude by observing that Eq. (3.91) implies
$\Upsilon(X)$ $=\log\upsilon_{0}=\log\left(\frac{{\mathcal{M}}_{11}(X)-{\mathcal{M}}_{22}(X)\pm\sqrt{\Delta_{Q}(X)}}{2{\mathcal{M}}_{12}(X)}\right)\,,$ (3.110a)
$\varphi_{Q}(X)$ $=\frac{1}{Q}\log\Pi\upsilon=\frac{1}{Q}\log\left(\frac{{\rm tr}{\mathcal{M}}_{Q}(X)\pm\sqrt{\Delta_{Q}(X)}}{2}\right)\,,$ (3.110b)

### v2

@@ 3.3 The wave function in the rational case, I
$\widetilde{\varphi}^{(+)}_{k}(X_{D})$ $=\sum_{s\geq 0}{\hat{a}_{sP+k}(q_{D})\over(sP+k)q_{D}^{\sigma_{sP+k}}}X_{D}^{sP+k},\qquad k=1,\dots,P-1,$ (3.40c)
$\widetilde{\varphi}^{(-)}_{k}(X_{D})$ $=\sum_{s\geq 0}{\hat{a}_{-(sP+k)}(q_{D})\over-(sP+k)q_{D}^{sP+k+\sigma_{-(sP+k)}}}X_{D}^{-(sP+k)},\qquad k=1,\dots,P-1,$ (3.40d)
@@
$\widetilde{\varphi}_{k}(X_{D})=\widetilde{\varphi}^{(+)}_{k}(X_{D})-q_{D}^{k}\widetilde{\varphi}^{(-)}_{k}(X_{D}),\qquad k=1,\dots,Q-1.$ (3.41b)
Putting equations (3.38) and (3.39) together and suitably identifying the functions just introduced in equation (3.41), we can now write the resummed series in the antiholomorphic block as
@@
$\widetilde{\varphi}^{(+)}_{P}(X_{D})={Q\over P}\varphi^{(+)}_{Q}(X),\qquad\widetilde{\varphi}^{(-)}_{P}(X_{D})={Q\over P}q^{sQ\left(1-{{Q\over P}}\right)}\varphi^{(-)}_{Q}(X),$ (3.43)
as a direct consequence of the cancellation symmetry in equation (3.23). Therefore, after summing equations (3.35) and (3.42) and using the relations in equation (3.43), we obtain the expression
$\phi^{\rm WKB}(X;q)+\phi^{\rm WKB}(X_{D};q_{D})$
$\qquad=\left(-{1\over 2}-{Q\over 2P}-{{\rm i}Q\over 2\pi P}\log X\right)\varphi_{Q}(X)+\left({{\rm i}Q\over 2\pi P}+2q{\partial\over\partial q}\right)\int_{0}^{X}\varphi_{Q}(X^{\prime}){{\rm d}X^{\prime}\over X^{\prime}}$
$\phantom{\qquad=}{}+sQ\left(1-{{Q\over P}}\right)\int_{0}^{X}\varphi^{(-)}_{Q}(X^{\prime}){{\rm d}X^{\prime}\over X^{\prime}}+\sum_{k=1}^{Q-1}{\varphi_{k}(X)\over q^{k}-1}+\sum_{k=1}^{P-1}{\widetilde{\varphi}_{k}\bigl(X^{Q\over P}\bigr)\over q^{k{Q^{2}\over P^{2}}}-1}.$ (3.44)
Besides, the power series newly defined in equations (3.34) and (3.41), which dictate the formula for the exact wave function at rational points by means of equation (3.44), are explicitly and uniquely determined by the function
@@
$\varphi^{(\pm)}_{k}(X)={1\over Q}\sum_{m=0}^{Q-1}q^{\mp mk}\Upsilon^{(\pm)}(q^{m}X,q),\qquad k=1,\dots,Q-1,$ (3.46b)
@@
$\widetilde{\varphi}^{(\pm)}_{k}(X_{D})={1\over P}\sum_{m=0}^{P-1}q_{D}^{\mp mk}\Upsilon^{(\pm)}(q_{D}^{m}X_{D}),\qquad k=1,\dots,P-1.$ (3.47b)
Let us stress that the derivation we have presented here implies a fixed choice of the classical branch. Indeed, the functions $\Upsilon(X)$, $\varphi_{Q}(X)$, $\varphi_{k}(X)$, $\widetilde{\varphi}_{P}(X_{D})$, $\widetilde{\varphi}_{k}(X_{D})$ are implicitly dependent on the label $\alpha$. We note that, applying equations (3.27) and (3.28), we can simply read off the integer coefficients $c_{\pm s}$, $s\in{\mathbb{Z}}_{>0}$, from the function $\varphi_{Q}(X)$. Namely, we have that
@@
after substituting $q^{sQ}=1$. In Appendix B, we describe an alternative way of computing the integer sequences $\{c_{\pm s}\}$, $s\in{\mathbb{Z}}_{>0}$ using equation (3.44). We will show in Section 4 that, starting from the known closed expression for the holomorphic and antiholomorphic blocks, we can successfully compute the integer coefficients for both series in $X$, $X^{-1}$ and all choices of the flat connection $\alpha$ in the examples of the figure-eight and three-twist knots.
@@ 3.4 The wave function in the rational case, II
In Section 3.3, we showed that the exact wave function $\chi(u;\hbar)$ in equation (3.13) at rational values of $\hbar$ is determined by the functions $\Upsilon^{(\alpha)}(X)$, defined in equation (3.45), and their derivatives with respect to $q$, up to the constants $\mathsf{C}_{\alpha}$ and the exponential prefactors containing $s^{(\alpha)}_{0,1}(x)$. We will now prove that the functions $\Upsilon^{(\alpha)}(X)$ can be obtained by solving directly the quantum $A$-polynomial in closed form. We will focus on $q$-difference equations of order two and three. These are the relevant cases for the examples studied in Section 4. The matrix formalism presented below was used in [44] for a second-order $q$-difference equation in the context of quantum mirror curves.
3.4.1 Solving a second-order $\bm{q}$-difference equation
Recall that, for each choice of $\alpha$, the resummed perturbative wave function $\chi^{(\alpha),{\rm WKB}}(u;\hbar)$ is given in equation (3.1). Equivalently, we write
$\chi^{(\alpha),{\rm WKB}}(u;\hbar)={\rm e}^{{{{\rm i}\over\hbar}}s^{(\alpha)}_{0}(x)+s^{(\alpha)}_{1}(x)}\Psi^{(\alpha),{\rm WKB}}(X;q),$ (3.49)
@@ 3.4.1 Solving a second-order $\bm{q}$-difference equation
$\Psi^{(\alpha),{\rm WKB}}(X;q)=\exp\bigl(\phi^{(\alpha),{\rm WKB}}(X;q)\bigr),$ (3.50)
@@
$\hat{A}\bigl(\hat{P},\hat{X},q\bigr)\chi^{(\alpha),{\rm WKB}}(u;\hbar)=0,$ (3.51)
is of second order. This is the case of the $\bm{4}_{1}$-knot, as shown in equation (2.30). After appropriately accounting for the correction due to the exponential prefactor containing the functions $s^{(\alpha)}_{0,1}(x)$ in equation (3.49), the $q$-difference equation above becomes
$\Psi^{(\alpha),{\rm WKB}}(qX;q)+A(X)\Psi^{(\alpha),{\rm WKB}}(X;q)+B(X)\Psi^{(\alpha),{\rm WKB}}\bigl(q^{-1}X;q\bigr)=0,$ (3.52)
where $A(X)$, $B(X)$ are functions of $X$ and $q$. Let us observe that, due to the conjectural structure in equations (3.2) and (3.3), we have that
$\upsilon^{(\alpha)}(X)={\Psi^{(\alpha),{\rm WKB}}(qX;q)\over\Psi^{(\alpha),{\rm WKB}}(X;q)}=\exp\bigl(\Upsilon^{(\alpha)}(X)\bigr),$ (3.53)
and we can now write equation (3.52) in terms of the above function $\upsilon^{(\alpha)}(X)$ as
$\upsilon(X)\upsilon\bigl(q^{-1}X\bigr)+A(X)\upsilon\bigl(q^{-1}X\bigr)+B(X)=0,$ (3.54)
where we have dropped the label $\alpha$ from the notation for simplicity. Let us further assume that $\hbar$ is fixed as in equation (3.15), that is,
$q={\rm e}^{{\rm i}\hbar}={\rm e}^{2\pi{\rm i}{P\over Q}},\qquad P,Q\in{\mathbb{Z}}_{>0}\ {\rm coprime},$ (3.55)
@@
$\upsilon_{k}=\upsilon\bigl(q^{-k}X\bigr).$ (3.56)
@@
$\upsilon_{k}\upsilon_{k+1}+A\bigl(q^{-k}X\bigr)\upsilon_{k+1}+B\bigl(q^{-k}X\bigr)=0,\qquad 0\leq k\leq Q-1,$ (3.57)
as a consequence of equation (3.54). We solve this system recursively, as follows. We define the sequences $\big\{a^{(k)}(X)\big\}$ and $\big\{b^{(k)}(X)\big\}$, $k\in{\mathbb{Z}}_{\geq 0}$, via the relations
@@
$\upsilon_{0}\upsilon_{1}\cdots\upsilon_{k}+a^{(k)}(X)\upsilon_{k}+b^{(k)}(X)=0,\qquad k>1.$ (3.58c)
If we multiply equation (3.58c) by $\upsilon_{k+1}$ and apply equation (3.57), we obtain a recursion relation which can be written in matrix form as
$\begin{pmatrix}a^{(k+1)}(X)\\ b^{(k+1)}(X)\end{pmatrix}=\begin{pmatrix}-A\bigl(q^{-k}X\bigr)&1\\ -B\bigl(q^{-k}X\bigr)&0\end{pmatrix}\begin{pmatrix}a^{(k)}(X)\\ b^{(k)}(X)\end{pmatrix}$, (3.59)
@@
${\cal M}_{Q}(X)=\prod_{k=1}^{Q}\begin{pmatrix}-A\bigl(q^{k}X\bigr)&1\\ -B\bigl(q^{k}X\bigr)&0\end{pmatrix}=\begin{pmatrix}{\cal M}_{11}(X)&{\cal M}_{12}(X)\\ {\cal M}_{21}(X)&{\cal M}_{22}(X)\end{pmatrix}$, (3.61)
@@
${\mathcal{M}}_{Q}\bigl(q^{-1}X\bigr)$ $=\begin{pmatrix}-A(X)&1\\ -B(X)&0\end{pmatrix}{\mathcal{M}}_{Q}(X)\begin{pmatrix}-A(X)&1\\ -B(X)&0\end{pmatrix}^{-1}$
$=\begin{pmatrix}{\cal M}_{22}-A{\cal M}_{12}&{{A\over B}}({\cal M}_{11}+A{\cal M}_{12})-{{1\over B}}({\cal M}_{21}+A{\cal M}_{22})\\ -B{\cal M}_{12}&{\cal M}_{11}+A{\cal M}_{12}\end{pmatrix}$, (3.62)
where we have removed the explicit dependence on $X$ for compactness. Following equations (3.60) and (3.62), we find that [^11] There is a sign misprint in the corresponding equation in [44].
$\begin{pmatrix}a^{(Q)}\bigl(q^{-1}X\bigr)\\ b^{(Q)}\bigl(q^{-1}X\bigr)\end{pmatrix}=\begin{pmatrix}{-\mathcal{M}}_{22}(X)+A(X){\mathcal{M}}_{12}(X)\\ B(X){\mathcal{M}}_{12}(X)\end{pmatrix}$. (3.63)
@@
which is invariant under $q$-shifts. We consider equation (3.58c) for $k=Q$ and, shifting $X$ into $q^{-1}X$, we produce the system
$\upsilon_{0}\bigl(\Pi\upsilon+a^{(Q)}(X)\bigr)+b^{(Q)}(X)=0,\qquad\upsilon_{1}\bigl(\Pi\upsilon+a^{(Q)}\bigl(q^{-1}X\bigr)\bigr)+b^{(Q)}\bigl(q^{-1}X\bigr)=0.$ (3.65)
By substituting the expressions for $a^{(Q)}(X)$, $a^{(Q)}\bigl(q^{-1}X\bigr)$ and $b^{(Q)}(X)$, $b^{(Q)}\bigl(q^{-1}X\bigr)$ in terms of the entries of the original matrix $\mathcal{M}_{Q}(X)$ in equations (3.60) and (3.63) and then using equation (3.57) with $k=0$, we can rewrite the system above as
$\upsilon_{0}(\Pi\upsilon-{\mathcal{M}}_{11}(X))-{\mathcal{M}}_{21}(X)=0,\qquad\Pi\upsilon-{\mathcal{M}}_{22}(X)-\upsilon_{0}{\mathcal{M}}_{12}(X)=0,$ (3.66)
@@
$\Pi\upsilon={{\operatorname{tr}{\mathcal{M}_{Q}}(X)\pm\sqrt{\Delta_{Q}(X)}\over 2}},$ (3.68)
@@
$\Delta_{Q}(X)=(\operatorname{tr}{\mathcal{M}_{Q}}(X))^{2}-4\det{\mathcal{M}_{Q}}(X),$ (3.69)
while $\det{\mathcal{M}_{Q}}(X)$ and $\operatorname{tr}{\mathcal{M}_{Q}}(X)$ denote the trace and determinant of the matrix, respectively. Note that, following equation (3.61), both $\det{\mathcal{M}_{Q}}(X)$ and $\operatorname{tr}{\mathcal{M}_{Q}}(X)$ are invariant under $q$-shifts and so depend on $X$ through $X^{Q}$. The function $\upsilon_{0}$ can be found, for example, using the second line of equation (3.66), yielding
$\upsilon_{0}={{\Pi v-{\mathcal{M}}_{22}(X)\over{\mathcal{M}}_{12}(X)}}={{{\mathcal{M}}_{11}(X)-{\mathcal{M}}_{22}(X)\pm\sqrt{\Delta_{Q}(X)}\over 2{\mathcal{M}}_{12}(X)}},$ (3.70)
while the functions $\upsilon_{k}$, $k\in{\mathbb{Z}}_{>0}$, are obtained by successively $q$-shifting $\upsilon_{0}$. This gives the solution to equation (3.54) for all rational values of $\hbar$. However, we stress that, for such a solution to be consistent with the original choice of a classical branch labeled by $\alpha$, we must impose that, taking $P=Q=1$, we find
@@
where $P^{(\alpha)}$ is the selected non-abelian solution of the classical $A$-polynomial at fixed $X$ after taking into account the corrections due to the exponential prefactor in equation (3.49). In this way, we remove the sign ambiguity in equation (3.68) and correspondingly in equation (3.70). We conclude by observing that equation (3.53) implies
$\Upsilon(X)$ $=\log\upsilon_{0}=\log\left({{{\mathcal{M}}_{11}(X)-{\mathcal{M}}_{22}(X)\pm\sqrt{\Delta_{Q}(X)}\over 2{\mathcal{M}}_{12}(X)}}\right),$ (3.72a)
$\varphi_{Q}(X)$ $={{1\over Q}}\log\Pi\upsilon={{1\over Q}}\log\left({{\operatorname{tr}{\mathcal{M}}_{Q}(X)\pm\sqrt{\Delta_{Q}(X)}\over 2}}\right),$ (3.72b)